# A Survey on Transformer Compression

### 05 Feb 2024 v1

@@ 4.2 Efficient Architecture
BERT [91] Encoder 0.1Bsimilar-$\sim$0.3B 12$\sim$24 12$\sim$16 3072$\sim$4096
PaLM [35] Decoder 9Bsimilar-$\sim$540B 32$\sim$118 16$\sim$48 4096$\sim$18432
Gopher [178] Decoder 4Msimilar-$\sim$280B 8$\sim$80 16$\sim$128 512$\sim$16384
BLOOM [147] Decoder 0.6Bsimilar-$\sim$176B 24$\sim$70 16$\sim$112 1024$\sim$14336
@@
ViT [127] Encoder 86Msimilar-$\sim$0.6B 12$\sim$32 12$\sim$16 3072$\sim$5120
DeiT [126] Encoder 5Msimilar-$\sim$86M 12 3$\sim$12 768$\sim$3072
TNT [182] Encoder 6Msimilar-$\sim$66M 12 3$\sim$10 768$\sim$2560
PVT [183] Pyramid 13Msimilar-$\sim$61M 12$\sim$45 8 2048
CMT [184] Pyramid 10Msimilar-$\sim$75M 24$\sim$40 8 1324$\sim$2432
Swin [38] Pyramid 29Msimilar-$\sim$0.2B 16$\sim$28 24$\sim$48 3072$\sim$6144
4.2.1 Representative Transformer-Based Models
Table VII presents the model cards of several representative LLMs and LVMs with public details. Vanilla transformer [11] is based on the encoder-decoder architecture [132]. In this architecture, the encoder stacks multi-head self-attention layers to encode the input sequence into latent representations, while the decoder performs cross-attention on these representations and generates target sequences in an autoregressive manner. Bidirectional Encoder Representations from Transformers (BERT) [91] is designed to pre-train deep bidirectional representations from unlabeled text by considering both left and right context in all layers. GPT-series models [4, 2] have effectively demonstrated the power of in-context learning using decoder-only architectures [147, 178, 181]. These decoder architectures incorporate unidirectional attention masks to ensure that each input token can only attend to previous tokens, including itself. The input and output tokens are processed in a similar manner within the decoder. In the following, we will provide separate reviews from the perspectives of CV and NLP on the current structural innovations in transformer-based architectures.
4.2.2 Language Domain
The self-attention mechanism in conventional transformer architectures often faces quadratic computational complexity. This poses a challenge for training and making inferences with long input sequences. To enhance efficiency, current innovations in the mainstream can be classified into three categories: optimizations targeting the attention mechanism, the direct replacement of attention with more efficient architectures, and enhancements focused on the Feed-Forward Network (FFN) component.
Attention Innovation. Various efficient Transformer variants have been proposed to reduce the computational complexity of the attention mechanism by incorporating structural priors on attention, such as sparsity [185, 186, 187, 188, 189, 190]. For example, Reformer [37] approximated full attention computation using locality-sensitive hashing, reducing the complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$. Other techniques include pooling-based compression [189], clustering methods [188, 191] that apply k-means clustering to learn dynamic sparse attention regions, and Longformer [192], which combines local windowed attention with task-motivated global attention. Chelba et al. [193] proposed truncating the target-side window based on an $N$-gram assumption, restricting the self-attention mechanism to use only the previous $N-1$ tokens. Additionally, locally banded sparse attention methods, such as Factorized Attention [187], have been adopted in models like GPT-3 [4].
Other methods have been proposed to reduce the computational complexity of the attention mechanism using linear attention. Performer [194] introduced fast attention via positive orthogonal random features, which is a scalable kernel method for approximating softmax attention kernels. This approach can efficiently model kernelizable attention mechanisms beyond softmax and comes with strong theoretical guarantees, including nearly unbiased estimation of the attention matrix, uniform convergence, and low estimation variance. Linear Transformer [195] reformulates self-attention as a linear dot-product of kernel feature maps. It uses kernels $\frac{\phi(q_{i})\phi(k_{j})}{\sum_{n=1}^{|x|}\phi(q_{i})\phi(k_{n})}$ to replace the softmax function and takes advantage of the associativity property of matrix products to reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
Moreover, there have been studies [35, 196, 197] exploring the concept of multi-query attention, where different heads share the same linear transformation matrices for the keys and values, as proposed by Shazeer et al.[197]. This approach offers significant computational savings with only a minimal impact on model quality. To further strike a balance between multi-query attention and multi-head attention, GQA et al. [198] introduces to group attention heads. In this approach, heads that belong to the same group share the same transformation matrices.
@@ 4.2.2 Language Domain
FlashAttention [199] offered an innovative approach to optimizing the speed and memory utilization of attention modules on GPUs. It focuses on an Input/Output (IO)-aware perspective, making better use of the fast memory SRAM by dividing the input into blocks. This optimization technique has already been integrated into various platforms [200, 201, 202, 203, 204] developed for training LLMs.
@@
Non-Transformer Architecture. Recent research has focused on developing new architectures for language modeling, including parameterized state space models [205, 206, 207, 208], long convolutions [209], and the incorporation of recurrent models [210, 211, 36, 212], which come with strong inductive biases. In particular, Gu et al. [205] introduced the S4 model, based on a novel parameterization for the SSM [213, 214], which involves conditioning matrix A with a low-rank correction, making it stably diagonalizable. This reduces the SSM to the well-studied computation of a Cauchy kernel [215]. The GSS model [206] further extended the use of gating units to the state space model family, and it was observed that employing gating units leads to dimensionality reduction when performing FFT operations, which addresses a major bottleneck in training speed. Hyena [209] trained a recurrence of gating units and implicitly parametrized long convolutions, which serves as an attention-free drop-in replacement for the traditional transformer architecture. Ma et al. [212] introduced MEGA, a single-head gated attention mechanism enhanced with exponential moving average. This mechanism aims to integrate the inductive bias of position-aware local dependencies into the position-agnostic attention mechanism.
@@
Peng et al. [216] proposed RWKV based on the fact that Recurrent Neural Networks (RNNs) show linear scaling in memory and computational requirements but encounter limitations in parallelization and scalability. They leveraged a linear attention mechanism, specifically replacing the quadratic QK attention with a scalar formulation that has linear cost. Additionally, they redefined recurrence and sequential inductive biases to parallelize computations during training while maintaining constant computational and memory complexity during inference.
@@
Furthermore, Retentive Network (RetNet) [36] theoretically derived the connection between recurrence and attention and proposed the retention mechanism for sequence modeling. This mechanism supports three computation paradigms, namely parallel, recurrent, and chunkwise recurrent, achieving training parallelism, low-cost inference, and good performance simultaneously. The parallel representation (also depicted in Figure 5(a)) of retention is:
@@
FFN Innovation. Large and sparse Feed-Forward Networks (FFNs) such as Mixture-of-Experts (MoE) [217, 218, 219, 220, 221, 222] have been effective in scaling up transformer-based models for pre-training LLMs. They replace a single FFN module with multiple equally-sized modules (experts) and activate only a few experts based on the input. This selective activation of FFN improves generalization performance while maintaining fixed training and inference costs. Liu et al. [223] further found a simpler selection method known as Avg-K, which selects blocks based on their mean aggregated hidden states. This method achieved lower perplexity in language model pre-training compared to existing MoE architectures such as Switch Transformer [219] and HashLayer [220].
Romero et al. [224] presented Continuous Kernel Convolution (CKConv) for handling arbitrarily long sequences in a parallel manner within a single operation. CKConv formulates convolutional kernels as vector-valued continuous functions parameterized by a small MLP instead of a sequence of independent weights. This MLP takes a time-step as input and generates the value of the convolutional kernel at the corresponding position, allowing the generation of kernels at different resolutions and arbitrarily sizes.
PaLM [35] and LLaMA [1] both utilized the SwiGLU Activation (Swish($xW$) $\dot{x}V$) for original FFN intermediate activations. This choice was based on the observation that SwiGLU activations have been shown to substantially enhance quality in compute-equivalent experiments [225] when compared to standard activation functions such as ReLU, GeLU, or Swish [225]. It’s important to note that using SwiGLU does require three matrix multiplications in the MLP instead of two.
@@ 4.2.3 Vision Domain
Dosovitskiy et al. [127] first introduced the Vision Transformer (ViT) primarily designed for image classification tasks. In the ViT model, the input image is initially divided into a sequence of fixed-length tokens, which are then processed through multiple transformer layers to capture global relationships among the tokens. Similar to the challenges faced in the language domain, the self-attention mechanism with quadratic computational complexity is one of the major hurdles for training and inference in transformer-based architectures applied to vision tasks. Many dense prediction tasks, such as object detection and image segmentation, often require higher-resolution inputs, necessitating the manual design of efficient architectures to handle these challenges. To address these issues, novel architectures like PVT [183] and Swin [183] have been proposed. These architectures incorporate hierarchical (pyramid) structures with multiple stages, effectively overcoming the challenges of adapting the original isotropic ViT [127] to various dense prediction tasks in computer vision.
@@
Enhancing Locality. T2T-ViT [242] leveraged token transformations to reduce token length by iteratively aggregating neighboring tokens to one token. LeViT [243] and MobileViT [244] employed hybrid architectures with stacked convolution layers, efficiently reducing the number of features through the first layer. TNT [182] further divided original ViT’s 16$\times$16 patches into smaller 4$\times$4 patches for capturing local information. CrossViT [231] processed small-patch and large-patch tokens separately, fusing them through multiple attention mechanisms. Twins [245] alternated between local and global attention layers for improved performance. RegionViT [246] introduced regional tokens and local tokens, enhancing local context with global information. KVT [247] introduced the k-NN attention to utilize locality of images patches and ignore noisy tokens by only computing attentions with top-k similar tokens. CMT [184] applied depth-wise convolutions to augment local patterns in the attention map and intermediate activation of the FFN. Pan et al. [248] proposed HiLo attention to disentangle the high/low frequency patterns in an attention layer by separating the heads into two groups, each equipped with specialized operations that focus on local windows.
@@
Faster Attention. Parmar et al. [233] introduced restrictions on the attention mechanism to focus on local neighborhoods. Swin [38] and Cswin [249] incorporated local attention within a window and introduced a shifted window partitioning method to enable cross-window connections. Shuffle Transformer [250] and Msg-transformer [251] employed spatial shuffle operations as alternatives to shifted window partitioning, facilitating cross-window connections. FasterViT [234] introduced hierarchical attention, breaking down global self-attention into multi-level attention components. This approach combines local window attention and hierarchical attention to achieve global information propagation while reducing computational costs. FLatten Transformer [252] integrated depth-wise convolution in conjunction with linear attention mechanisms to address the challenge of maintaining diversity in output features across different positions.
@@
Attention-Free Architecture. AFT [253] pioneered an innovative approach by combining key and value elements with learned position biases, followed by element-wise multiplication with the query. This operation’s distinctive feature is its linear memory complexity concerning both context size and feature dimension, making it compatible with large input and model sizes. GFnet [254] presented an alternative by replacing traditional attention mechanisms with Fast Fourier Transform (FFT), frequency gating, and inverse FFT for rapid token mixing. Additionally, some architectures have been entirely based on pure multi-layer perceptrons (MLPs) [255, 40, 256, 257, 240, 241], without using convolutions or self-attention mechanisms. Yu et al. [39] proposed MetaFormer as a general architecture abstracted from ViTs without specifying the token mixer. By employing basic token mixing, primarily through non-parametric pooling, MetaFormer achieved satisfactory performance.
@@ 4.2.4 Discussion
Transformer models consistently achieve state-of-the-art results across various tasks, but the cost of training these models, especially on long sequences, can be prohibitive. Table VIII concludes several papers that propose improvements to the attention mechanism from three different dimensions. While using linear attention can help reduce the computational complexity of the attention mechanism, it struggles to effectively encode position information, leading to a potential decrease in model performance. Models like RWKV [216] and RetNet [36], which generate outputs recursively similar to RNNs, present a promising direction for further exploration. These models only need to refer to the previous state during decoding, significantly improving decoding efficiency by avoiding the need to revisit all previous states, a characteristic of conventional Transformers. Additionally, they can encode entire sentences in parallel, enabling highly parallel and efficient training.
@@ 5 Other Approaches
In addition to quantization, distillation, pruning, and novel network architectures, there are also several other model compression and acceleration approaches.
@@
Tensor Decomposition. Tensor or matrix decomposition aims to decompose the large tensor or matrix into smaller ones so as to save the number of parameters and computational costs. This approach is firstly introduced into compression of fully-connected layers and convolutional networks [263, 264]. As for large language models, Tensor decomposition is utilized for simplifying model weight or embedding layers. Edalati et.al [265] is the first attempt to use Kronecker decomposition for compressing generative language models. This work represent weight matrix with two smaller matrices using Kronecker decomposition and achieves lossless 1.5×1.5$\times$1.5 × compression of GPT2. LoRD [266] further use low rank decomposition to compress code LLMs. TSVD [267] compress linear mappings in transformers using SVD and constrain $U$ and $V$ matrices in ternary format. In addition to model weights, TensorGPT [268] proposes to compress embedding layers and achieves more than 3×3$\times$3 × compression factor with lossless performance.
@@
Early Exiting. Early exiting can dynamically allocate different resources for each input sample and maintain the original performance, which has been successfully used in information retrieval system [269] and convolutional networks [270]. Many early exiting technologies have been proposed for encoder-only Transformers [271, 272, 273, 274]. The key problem in early exiting is determining when to exit. The existing works mainly utilize intrinsic confidence measures [271], routing in advance [275] or training a early-exit classifier [276]. For transformers with decoders, several works have been proposed. Elbayad et.al [277] introduces the depth-adaptive transoformer for accelerating machie translation. CALM [278] proposes to dynamically allocate different computational resources for different input and timestep. CALM can achieve speedup of up to 3×3$\times$3 × on T5 encoder-decoder model. SkipDecoder [279] bypasses tokens in lower layers to middle layers which enable batch inference and reuse KV caching. SkipDecoder obtains 2×2$\times$2 × to 5×5$\times$5 × inference speedups on OPT models.
@@
Speculative Sampling. Speculative sampling is a special accelerating approach for transformer decoding by computing several tokens in parallel [280, 281]. In large language model, decoding $K$ tokens requires $K$ runs of the model which is slow. Taking the reference tokens generated from smaller models, speculative sampling runs these tokens in parallel can significantly make the decoding process faster. Moreover, the rejection scheme [281] can preserve the distribution of the original LLM so that speculative sampling is theoretically lossless. Yang et.al [282] further take the input text as reference tokens without introducing extra models. LLMCad [283] introduces an on-device inference engine for LLMs with speculative sampling as a key technology.
@@
Discussion. We have reviewed the methods that are different from the above in this section, including tensor decompostion, early exiting and speculative sampling. These methods do not change the basic operations, rather accelerate model inference by modifing the network connections. Tensor decompostion approximates the original weights with low-rank ones which may have performance degradation. Early exiting makes prediciton by adaptively allocate resources per sample. Speculative sampling considers the specialization of decoders and can accelerate decoding losslessly. These special methods are inspiring for developing new algorithms for accelerating LLMs or LVMs in the future.

### 07 Apr 2024 v2

@@ 4.2.2 Language Domain
BERT [91] Encoder 0.1B-0.3B 12-24 12-16 3072-4096
PaLM [35] Decoder 9B-540B 32-118 16-48 4096-18432
Gopher [179] Decoder 4M-280B 8-80 16-128 512-16384
BLOOM [147] Decoder 0.6B-176B 24-70 16-112 1024-14336
@@
ViT [128] Encoder 86M-0.6B 12-32 12-16 3072-5120
DeiT [127] Encoder 5M-86M 12 3-12 768-3072
TNT [197] Encoder 6M-66M 12 3-10 768-2560
PVT [198] Pyramid 13M-61M 12-45 8 2048
CMT [199] Pyramid 10M-75M 24-40 8 1324-2432
Swin [38] Pyramid 29M-0.2B 16-28 24-48 3072-6144
@@
FlashAttention [200] offers an innovative approach to optimizing the speed and memory utilization of attention modules on GPUs. It focuses on an Input/Output (IO)-aware perspective, making better use of the fast memory SRAM by dividing the input into blocks. This optimization technique has already been integrated into various platforms [201, 202, 203, 204, 205] developed for training LLMs.
@@
Non-Transformer architecture. Recent research has focused on developing new architectures for language modeling, including parameterized state space models [206, 207, 208, 209], long convolutions [210], and the incorporation of recurrent models [211, 212, 36, 213], which come with strong inductive biases. In particular, Gu et al. [206] introduces the S4 model, based on a novel parameterization for the SSM [214, 215], which involves conditioning matrix A with a low-rank correction, making it stably diagonalizable. This reduces the SSM to the well-studied computation of a Cauchy kernel [216]. The GSS model [207] further extends the use of gating units to the state space model family, and it is observed that employing gating units leads to dimensionality reduction when performing FFT operations, which addresses a major bottleneck in training speed. Hyena [210] trains a recurrence of gating units and implicitly parametrized long convolutions, which serves as an attention-free drop-in replacement for the traditional Transformer architecture. Ma et al. [213] introduces MEGA, a single-head gated attention mechanism enhanced with exponential moving average. This mechanism aims to integrate the inductive bias of position-aware local dependencies into the position-agnostic attention mechanism.
@@
Peng et al. [217] proposes RWKV based on the fact that Recurrent Neural Networks (RNNs) show linear scaling in memory and computational requirements but encounter limitations in parallelization and scalability. They leverage a linear attention mechanism, specifically replacing the quadratic QK attention with a scalar formulation that has linear cost. Additionally, they redefine recurrence and sequential inductive biases to parallelize computations during training while maintaining constant computational and memory complexity during inference.
@@
Furthermore, Retentive Network (RetNet) [36] theoretically derives the connection between recurrence and attention and proposes the retention mechanism for sequence modeling. This mechanism supports three computation paradigms, namely parallel, recurrent, and chunkwise recurrent, achieving training parallelism, low-cost inference, and good performance simultaneously. The parallel representation (also depicted in Figure 5(a)) of retention is:
@@
More recently, Gu et al. [178] have highlighted the limitations of previous structured SSMs in effectively modeling discrete and information-dense data, such as text. In response, they propose Mamba, consisting a novel class of selective mechanisms and hardware-aware designs that enable linear scalability to billions of parameters, while maintaining computational efficiency and achieving strong performance. Specifically, in comparison to the H3 [208] block, the Mamba model replaces the first multiplicative gate with an activation function. Additionally, Mamba incorporates an SSM into the main branch, distinguishing it from the MLP block in transformer. Building upon the foundation of Mamba, researchers are further unlocking the potential of SSMs for scaling by exploring new techniques, such as combining SSMs with Mixture of Experts (MoE) [218] or introducing methods to enhance the flow of hidden information [219] between layers in SSMs.
FFN innovation. Large and sparse FFNs such as Mixture-of-Experts (MoE) [220, 221, 222, 223, 224, 225] have been effective in scaling up Transformer-based models for pre-training LLMs. They replace a single FFN module with multiple equally-sized modules (experts) and activate only a few experts based on the input. This selective activation of FFN improves generalization performance while maintaining fixed training and inference costs. Liu et al. [226] further finds a simpler selection method known as Avg-K, which selects blocks based on their mean aggregated hidden states. This method achieves lower perplexity in language model pre-training compared to existing MoE architectures such as Switch Transformer [222] and HashLayer [223].
Romero et al. [227] presents Continuous Kernel Convolution (CKConv) for handling arbitrarily long sequences in a parallel manner within a single operation. CKConv formulates convolutional kernels as vector-valued continuous functions parameterized by a small MLP instead of a sequence of independent weights. This MLP takes a time-step as input and generates the value of the convolutional kernel at the corresponding position, allowing the generation of kernels at different resolutions and arbitrarily sizes.
PaLM [35] and LLaMA [1] both utilize the SwiGLU Activation (Swish($xW$) $\dot{x}V$) for original FFN intermediate activations. This choice is based on the observation that SwiGLU activations have been shown to substantially enhance quality in compute-equivalent experiments [228] when compared to standard activation functions such as ReLU, GeLU, or Swish [228]. It’s important to note that using SwiGLU does require three matrix multiplications in the MLP instead of two.
@@ 4.2.3 Vision Domain
Dosovitskiy et al. [128] first introduces the Vision Transformer (ViT) primarily designed for image classification tasks. In the ViT model, the input image is initially divided into a sequence of fixed-length tokens, which are then processed through multiple Transformer layers to capture global relationships among the tokens. Similar to the challenges faced in the language domain, the self-attention mechanism with quadratic computational complexity is one of the major hurdles for training and inference in Transformer-based architectures applied to vision tasks. Many dense prediction tasks, such as object detection and image segmentation, often require higher-resolution inputs, necessitating the manual design of efficient architectures to handle these challenges. To address these issues, novel architectures like PVT [198] and Swin [198] have been proposed. These architectures incorporate hierarchical (pyramid) structures with multiple stages, effectively overcoming the challenges of adapting the original isotropic ViT [128] to various dense prediction tasks in computer vision.
@@
Enhancing locality. T2T-ViT [245] leverages token transformations to reduce token length by iteratively aggregating neighboring tokens to one token. LeViT [246] and MobileViT [247] employ hybrid architectures with stacked convolution layers, efficiently reducing the number of features through the first layer. TNT [197] further divides original ViT’s 16$\times$16 patches into smaller 4$\times$4 patches for capturing local information. CrossViT [234] processes small-patch and large-patch tokens separately, fusing them through multiple attention mechanisms. Twins [248] alternates between local and global attention layers for improved performance. RegionViT [249] introduces regional tokens and local tokens, enhancing local context with global information. KVT [250] introduces the k-NN attention to utilize locality of images patches and ignore noisy tokens by only computing attentions with top-k similar tokens. CMT [199] applies depth-wise convolutions to augment local patterns in the attention map and intermediate activation of the FFN. Pan et al. [251] proposes HiLo attention to disentangle the high/low frequency patterns in an attention layer by separating the heads into two groups, each equipped with specialized operations that focus on local windows.
@@
Faster attention. Parmar et al. [236] introduces restrictions on the attention mechanism to focus on local neighborhoods. Swin [38] and Cswin [252] incorporate local attention within a window and introduce a shifted window partitioning method to enable cross-window connections. Shuffle Transformer [253] and Msg-Transformer [254] employ spatial shuffle operations as alternatives to shifted window partitioning, facilitating cross-window connections. FasterViT [237] introduces hierarchical attention, breaking down global self-attention into multi-level attention components. This approach combines local window attention and hierarchical attention to achieve global information propagation while reducing computational costs. FLatten Transformer [255] integrates depth-wise convolution in conjunction with linear attention mechanisms to address the challenge of maintaining diversity in output features across different positions.
@@
Attention-free architecture. AFT [256] pioneers an innovative approach by combining key and value elements with learned position biases, followed by element-wise multiplication with the query. This operation’s distinctive feature is its linear memory complexity concerning both context size and feature dimension, making it compatible with large input and model sizes. GFnet [257] presents an alternative by replacing traditional attention mechanisms with Fast Fourier Transform (FFT), frequency gating, and inverse FFT for rapid token mixing. Additionally, some architectures have been entirely based on pure multi-layer perceptrons (MLPs) [258, 40, 259, 260, 243, 244], without using convolutions or self-attention mechanisms. Yu et al. [39] proposes MetaFormer as a general architecture abstracted from ViTs without specifying the token mixer. By employing basic token mixing, primarily through non-parametric pooling, MetaFormer achieves satisfactory performance.
Inspired by the efficient hardware-aware designs of Mamba in NLP, Zhu et al. [261] propose a novel generic vision backbone called Vim with bidirectional Mamba blocks. Different from the sequential nature of language, representing visual data in SSMs presents challenges primarily due to their sensitivity to spatial position. To address this, Vim introduces position embeddings to mark image sequences and compresses visual representations using bidirectional SSM. Following the architectures of ViT [12] and DeiT [26], Vim initially employs a 16×16 kernel size projection layer to obtain a 1-D sequence of non-overlapping patch embeddings. Simultaneously, Liu et al. [262] propose the visual state space model (VMamba), which introduces the Cross-Scan Module (CSM) to traverse the spatial domain and convert any non-causal visual image into ordered patch sequences, thereby addressing the inherent direction-sensitive issue. Following the architecture of Swin [136], VMamba begins the process by partitioning the input image into patches using a stem module, without further flattening the patches into a 1-D sequence.
@@ 4.2.4 Discussion
Transformer models consistently achieve state-of-the-art results across various tasks, but the cost of training these models, especially on long sequences, can be prohibitive. Table VIII concludes several papers that propose improvements to the attention mechanism from three different dimensions. While using linear attention can help reduce the computational complexity of the attention mechanism, it struggles to effectively encode position information, leading to a potential decrease in model performance. Models like RWKV [217], RetNet [36] and Mamba [178], which generate outputs recursively similar to RNNs, present a promising direction for further exploration. These models only need to refer to the previous state during decoding, significantly improving decoding efficiency by avoiding the need to revisit all previous states, a characteristic of conventional Transformers. Additionally, they can encode entire sentences in parallel, enabling highly parallel and efficient training.
@@ 5 Specialized Approaches
In addition to quantization, distillation, pruning, and novel network architectures, there are several other model compression and acceleration approaches including tensor decomposition, early exiting, and speculative sampling.
@@
Tensor decomposition. Tensor or matrix decomposition aims to decompose a large tensor or matrix into smaller ones in order to reduce the number of parameters and computational costs. This approach was first introduced into the compression of fully-connected layers and convolutional networks [268, 269]. In terms of large language models, Tensor decomposition is utilized for simplifying model weight or embedding layers. Edalati et.al [270] is the first attempt to use Kronecker decomposition for compressing generative language models. This work represent weight matrix with two smaller matrices using Kronecker decomposition and achieves lossless 1.5×1.5$\times$1.5 × compression of GPT2. LoRD [271] further uses low rank decomposition to compress code LLMs. And TSVD [272] compresses linear mappings in Transformers using SVD and constrain $U$ and $V$ matrices in ternary format. In addition to model weights, TensorGPT [273] proposes to compress embedding layers and achieves more than 3×3$\times$3 × compression factor with lossless performance.
@@
Early exiting. Early exiting can dynamically allocate different resources for each input sample and maintain the original performance. This technology has been successfully used in information retrieval system [274] and convolutional networks [275]. Many early exiting technologies have been proposed for encoder-only Transformers [276, 277, 278, 279]. The key problem in early exiting is determining when to exit. The existing works mainly utilize intrinsic confidence measures [276], routing in advance [280], or training a early-exit classifier [281]. For Transformers with decoders, several works have been proposed. For example, Elbayad et.al [282] introduces the depth-adaptive Transformer for accelerating machine translation. CALM [283] proposes to dynamically allocate different computational resources for different input and timesteps, achieving speedups of up to 3×3$\times$3 × on T5 encoder-decoder model. Additionally, SkipDecoder [284] bypasses tokens in lower layers to middle layers in order to enable batch inference and reuse KV caching, achieving inference speedups of 2×2$\times$2 × to 5×5$\times$5 × on OPT models.
@@
Speculative sampling. Speculative sampling, a special acceleration approach for Transformer decoding, computes several tokens in parallel [285, 286]. In large language models, decoding $K$ tokens requires $K$ runs of the model – a slow process. Taking the reference tokens generated from smaller models, speculative sampling runs these tokens in parallel to significantly accelerate the decoding process. Moreover, the rejection scheme [286] can preserve the distribution of the original LLM so that speculative sampling is theoretically lossless. For example, Yang et.al [287] further take the input text as reference tokens without introducing extra models. And LLMCad [288] introduces an on-device inference engine for LLMs with speculative sampling as a key technology.
@@
Discussion. This section has reviewed the methods that differ from those mentioned earlier, covering tensor decomposition, early exiting, and speculative sampling. These methods do not change the basic operations but instead accelerate model inference by modifying the network connections. Tensor decomposition approximates the original weights with low-rank ones, potentially leading to performance degradation. Early exiting makes predictions by adaptively allocating resources on a per sample basis. And speculative sampling considers the specialization of decoders and can accelerate decoding in a lossless manner. These special methods inspire future development of new algorithms for accelerating LLMs or LVMs.